# AIRS-Bench: a Suite of Tasks for Frontier AI Research Science Agents

### v2

@@ 1 Introduction
From the rise of deep learning through to the current era of Large Language models (LLMs), the rapid progress in machine learning (ML) has been largely benchmark-driven, making robust evaluations a key requirement (Hardt, 2025). This trend is reinforced by the current reviewing crisis, which highlights the limits of human judgment in assessing the quality of research contributions (Xu et al., 2022; Lawrence, 2022). Recent improvements in LLM technology have extended their abilities from simple tasks to complex agentic workflows, including scientific reasoning and coding (Khatri et al., 2025; Andrews et al., 2025). Whilst early LLM research focused on improving the capabilities of LLMs through better data and pre-training, recent work has focused on extending their abilities by leveraging more test-time compute (Yao et al., 2023a), in particular by leveraging scaffolds that iteratively query the LLM at test-time so as to obtain a better response. Such scaffolds can enable a test-time search of the solution space by incorporating environment feedback (Nathani et al., 2025; Toledo et al., 2025) and yield much more performant solutions. Despite this potential, we lack a standardized framework to measure how well these agents perform the actual work of a research scientist. To fill this gap, we introduce AIRS-Bench.
@@ 6.1 Comparing performance across agents
Figure 4 provides an overview of the three aggregate benchmark metrics introduced above. We observe that reasoning models (e.g. gpt-oss-120b, o3-mini) perform better in both one-shot and greedy settings with model size also affecting performance, i.e. gpt-oss-120b outperforms gpt-oss-20b by a significant margin. Moreover, tree-search methods benefit agents powered by both open-source and closed-source models, as suggested by e.g. the sizeable gaps between the performances of Greedy CWM and One-Shot CWM as well as Greedy GPT-4o and One-Shot GPT-4o agents. At the same time, performance of agents backed by linear scaffolds, such as ReAct CWM and ReAct GPT4o, stands in the middle. We observe that the relative ranking of agents shows similar trends for all three performance metrics; the ability to submit a valid solution correlates with the ability to submit performant solutions and the Elo ranking. We also notice that models such as o3-mini demonstrate high participation but also high loss rates, suggesting a tendency to submit more frequently but with less selectivity; in contrast, models like CWM participate less often but with higher confidence, reflecting distinct agent “personalities” and risk profiles. Finally, the majority of agent-task combinations achieve results between the o3-mini baseline and human SOTA, with a small but notable fraction (1.55%) exceeding SOTA, primarily driven by greedy search strategies.
@@
We report the mean valid submission rate defined in Equation 1 across the AIRS-Bench tasks in Figure. 7. On average, only 59.3% of the total submissions are considered valid, suggesting that even submitting a valid solution stretches the capabilities of the agents. We also observe that reasoning models and the greedy scaffolds offer an advantage, as performance of these agents is superior.

### v3

@@ 1 Introduction
From the rise of deep learning through to the current era of Large Language models (LLMs), the rapid progress in machine learning (ML) has been largely benchmark-driven, making robust evaluations a key requirement (Hardt, 2025). This trend is reinforced by the current reviewing crisis, which highlights the limits of human judgment in assessing the quality of research contributions (Xu et al., 2022; Lawrence, 2022). Recent improvements in LLM technology have extended their abilities from simple tasks to complex agentic workflows, including scientific reasoning and coding (Khatri et al., 2025; Andrews et al., 2025; Lu et al., 2024; Yamada et al., 2025). Whilst early LLM research focused on improving the capabilities of LLMs through better data and pre-training, recent work has focused on extending their abilities by leveraging more test-time compute (Yao et al., 2023a), in particular by leveraging scaffolds that iteratively query the LLM at test-time so as to obtain a better response. Such scaffolds can enable a test-time search of the solution space by incorporating environment feedback (Nathani et al., 2025; Toledo et al., 2025) and yield much more performant solutions. Despite this potential, we lack a standardized framework to measure how well these agents perform the actual work of a research scientist. To fill this gap, we introduce AIRS-Bench.
@@ 6.1 Comparing performance across agents
Figure 4 provides an overview of the three aggregate benchmark metrics introduced above. We observe that reasoning models (e.g. gpt-oss-120b, o3-mini) perform better in both one-shot and greedy settings. However, model size seems to not affect performance with the use of test-time scaling making up for it, i.e. Greedy gpt-oss-120b is on par with Greedy gpt-oss-20b. Moreover, tree-search methods benefit agents powered by both open-source and closed-source models, as suggested by e.g. the sizeable gaps between the performances of Greedy CWM and One-Shot CWM as well as Greedy GPT-4o and One-Shot GPT-4o agents. At the same time, performance of agents backed by linear scaffolds, such as ReAct CWM and ReAct GPT4o varies; in one case the ReAct agent is better than its greedy counterpart (CWM) and in the other case worse (GPT4o). We observe that the relative ranking of agents shows similar (but not identical) trends for the average normalized score and Elo Rating performance metrics. However, the ability to submit a valid solution does not always correlate with the ability to submit a high-performing solution. This suggests that agents who are willing to take risks may achieve better performance. Finally, the majority of agent-task combinations achieve results between the worst possible score and human SOTA, with a small but notable fraction (1.58%) exceeding SOTA, primarily driven by greedy search strategies.
@@
We report the mean valid submission rate defined in Equation 1 across the AIRS-Bench tasks in Figure. 7. On average, only 55.1% of the total submissions are considered valid, suggesting that even submitting a valid solution stretches the capabilities of the agents. We also observe that reasoning models and the Greedy/ReAct scaffolds offer an advantage, as performance of these agents is superior.